William has scored a total of 235 points on the first three math tests taken in his mathematics class. If his mathematics tests are always worth 100 points, how many points must he earn on his next test in order to have an average of 80%?


a.) 80
b.) 85
c.) 90
d.) 95
Thank you in advance! Btw, if you can, could you please also show step by step examples on how to solve and do slopes on charts?

Answer by mananth(16949)

assume he has to get x points to get an average of 80

His total points will be 235+x
Number of tests will be 4
(235+x)/4 = 80
multiply by 4
235+x= 4*80
235+x= 320
x=320-235
x= 85
he has to get 85 points to average 80